OpenStudy (anonymous):

When an altitude is drawn to the hypotenuse of a right triangle, the lengths of the segments of the hypotenuse are 16 and 25. What is the length of the altitude?

OpenStudy (anonymous):

20 25 41 50 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

If you draw the picture, you'll have 3 right triangles. One with legs x and y, with hypotenuse 41 One with legs A and 25, with hypotenuse y and another one with legs 16 and A, with hypotenuse x So use the pythagorean theorem to get the three equations x^2+y^2 = 41^2 A^2+25^2 = y^2 16^2+A^2 = x^2 Then solve for A 16^2+A^2+y^2 = 41^2 16^2+A^2+A^2+25^2 = 1681 256+2A^2+625 = 1681 2A^2 = 1681 2A^2 = 800 A^2 = 400 A = 20 So the answer is choice A